OpenStudy (anonymous):

Find the absolute maximum and minimum of f(x, y) = 2x+3y subject to the constraint squareroot(x) + squareroot(y) = 5

OpenStudy (anonymous):

Use Lagrange multipliers

OpenStudy (anonymous):

Do you know how to find ReverseTriangle.F and ReverseTriangle.G?

OpenStudy (anonymous):

No i do not.

OpenStudy (anonymous):

Have you learned about Lagrange Multipliers lol?

OpenStudy (dan815):

how can u ask this questions then xD

OpenStudy (anonymous):

Use f(x, y) = 2x+3y to find ReverseTriangle.F

OpenStudy (anonymous):

Then use squareroot(x) + squareroot(y) = 5 to find ReverseTriangle.G

OpenStudy (dan815):

do u want to see a great video on lagrange multipliers?

OpenStudy (dan815):

very useful to know

OpenStudy (anonymous):

I have learned lagrange multipliers however I can only seem to find one critical point (9,4). The answer I have tells me that the minimum is (9.4) but the maximum is (0,25). How would you find the point (0,35)? I have not heard of this reverse triangle stuff

OpenStudy (dan815):

he just means delta F and delta G

OpenStudy (anonymous):

ohhh i see.

OpenStudy (anonymous):

well I still cannot find point (0,25)

OpenStudy (anonymous):

as a critical point.

OpenStudy (anonymous):

Find ReverseTriangle.F and ReverseTriangle.G first! Then we'll tell you how to go from there :)

OpenStudy (anonymous):

gradient f = <2, 3> gradient g = <1/(2squareroot(x)),1/(2squareroot(y))>

OpenStudy (anonymous):

Solving everything down gives me x=3/2root(y) which I plug in to g to give me x=9 Then pluging it back gives me the point (9.4)

OpenStudy (anonymous):

How do I get (0,25) which is the maximum?

OpenStudy (anonymous):

Ok so you have your two ReverseTriangles! Now we are getting somewhere. Now you want to make each component in your ReverseTriangle.F as a relationship with your ReverseTriangle.G times a Scalar

OpenStudy (anonymous):

\[\Delta F = \lambda \Delta G\]

OpenStudy (anonymous):

I have done that but solving the equation for x and substituting it into g gives me only the point (9,4)

OpenStudy (anonymous):

The key here is that I have no Idea how to get the point (0,25) as a critical point which the answer tells me is the maximum.

OpenStudy (anonymous):

Put in (9,4) back into function F!

OpenStudy (anonymous):

It gives me 30

OpenStudy (anonymous):

Are you sure you solved for lambda properly?

OpenStudy (anonymous):

lambda is equal to 12

OpenStudy (anonymous):

\[<2,3> = \lambda <\frac{ 1 }{ 2\sqrt{x} },\frac{ 1 }{ 2\sqrt{y} }>\]

OpenStudy (anonymous):

yes I have that.

OpenStudy (anonymous):

\[2=\frac{ \lambda }{ 2\sqrt{x} }, 3=\frac{ \lambda }{ 2\sqrt{y} }\]

OpenStudy (anonymous):

\[4\sqrt{x}=6\sqrt{y}\]

OpenStudy (anonymous):

Yes I have all of that. The one problem is that using all of this, I am unable to find the point (0,25) as a critical point when it clearly is one.

OpenStudy (anonymous):

What does Y equal in terms of X?

OpenStudy (anonymous):

(4/9)x

OpenStudy (anonymous):

squareroot(x) + squareroot(\[\sqrt{x}+\sqrt{\frac{ 4x }{ 9 }}=5\]) = 5

OpenStudy (anonymous):

\[=\sqrt{x}+\frac{ \sqrt{4x} }{ \sqrt{9} }=\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

\[\sqrt{x}+\frac{ 2\sqrt{x} }{ 3 }=5\]

OpenStudy (anonymous):

\[=3\sqrt{x}+2\sqrt{x}=15\]

OpenStudy (anonymous):

\[=5\sqrt{x}=15\]

OpenStudy (anonymous):

..lmao..

OpenStudy (anonymous):

..I see where this is heading..give me a sec

OpenStudy (anonymous):

I have all of this. The end result is x=9 and plugging that into (4/9)x=y gives y=4. The question is how I can find out (0,25) is a critical point.

OpenStudy (anonymous):

Is (25,0) also a critical point?

OpenStudy (anonymous):

According to the answers, (25,0) is the maximum while the (9,4) i've already found is the minimum.

OpenStudy (anonymous):

(0,25) sry

OpenStudy (anonymous):

I just need to know how to find out that (0,25) is a critical point.

OpenStudy (anonymous):

I'll be back. While I'm gone, let's see if The Big Guns, AKA @oldrin.bataku can help.

zepdrix (zepdrix):

Try to think back to Calc I. When we had a function and needed to find critical points, we would just use the derivative (very similar to what we're doing here with the Lagrange thingy-ma-bobber). But when we had a function `within an interval` we also had to check the end points of that interval. The same thing is happening here. So you've found critical points within the region, good. Now we need to check the boundaries of the region as well. We can do that by letting one of the variables be 0. \[\large x=0 \qquad \rightarrow \qquad \sqrt{y}=5 \qquad \rightarrow \qquad y=25\] Plugging these values into our original function gives us,\[\large f(0,25)=2\cdot0+3\cdot25\] It would also make sense to check the other boundary I guess, \[\large y=0\] Understand how that works?

OpenStudy (anonymous):

Ah... I had always assumed that lagrange would find the critical points on the boundaries as well. Thanks for your help!

OpenStudy (anonymous):

@FutureMathProfessor \(\nabla=\text{\nabla}\)

OpenStudy (anonymous):

\[\nabla\]

OpenStudy (anonymous):

W00T!